OpenStudy (anonymous):

Which of the following are equal to the expression below ? Check all that apply . sqrt9 / sqrt 81 a.1/3 b.2/9 c.3/9 d.3 ?? help someone plz thank you ????

OpenStudy (cwrw238):

sqrt9 / sqrt 81 = 3 /9 = 1 /3 a and c

OpenStudy (anonymous):

a and c